Given the following information for country A:

Details RM (Million)

Consumers consumption

30

Net investment

20

Government expenditure

30

Export

20

Import

10

Depreciation

10

Indirect taxes

25

Subsidies

118

a) Calculate Gross Domestic Product for country A. (4 marks)


1
Expert's answer
2021-08-29T16:37:40-0400

Gross Domestic Product can be determined at two levels:

(i) At factor cost:

GDP = C + I + G + Subsidies - Indirect Taxes

Consumers consumption (C) = 30

Net investment (I) = 20

Government expenditure (G) = 30


Indirect taxes = 25

Subsidies = 118


"\therefore" GDP = RM (30 + 20 + 30 + 118 - 25)

= RM. 173 million


(ii) Market Price;


GDP = C + I + G - Subsidies + Indirect Taxes

= 30 + 20 + 30 - 118 + 25

= RM -13million


NOTE: In the determination of Gross Domestic Product (GDP), Expenditure on Net Exports (Exports minus Imports or X - M) and Depreciation are not included.